# Quantifying Emotional Arousal through Pupillary Response: A Novel Approach for Isolating the Luminosity Effect and Predicting Affective States

### v1

@@ III-C Testing models developed by other researchers with our data
IV Discussion
In this study, we developed a novel mathematical model to remove the effects of stimulus and ambient luminosity and isolate emotional arousal information from measures of pupil size. This model serves as a simple yet effective and robust tracking system that can be used as a standalone physiological indicator for emotion arousal detection or can be integrated into more complex machine learning models. Our approach was designed to overcome longstanding methodological obstacles in the field, most notably the confounding effects of dynamic changes in luminosity on pupil dilation, which have historically obscured accurate emotion detection.
Prior studies in emotion detection have addressed the impact of luminosity on pupillary responses using methods such as Linear Regression [tarnowski2020eye, raiturkar2016decoupling], controlled visual stimuli [kuchinke2011spontaneous], isoluminant conditions [nakayama2021controlling], and baseline correction [bradley2008pupil]. For instance, LR models the relationship between pupil size and luminosity, enabling the subtraction of luminosity effects from pupil measurements [tarnowski2020eye]. Controlled visual stimuli and isoluminant conditions provide stable or uniform luminosity across experimental stimuli to minimize variability. Baseline correction uses pupil size measurements from a gray screen stimulus, subtracting them from emotion stimulus pupil sizes to remove the influence of luminosity and account for individual differences [bradley2008pupil]. Other more advanced methods further incorporate localized gaze luminosity alongside average luminosity to improve accuracy [bradley2008pupil, kuchinke2011spontaneous, nakayama2021controlling, raiturkar2016decoupling, partala2003pupil, holmqvist2011eye]. However, those methods have considerable limitations.
For example, Nakayama et al. [nakayama2021controlling] developed a method to predict pupil size as a function of computer screen luminosity in a darkroom. Here, we tested that method and obtained inferior results when applying their method to our stimulus set. The average coefficient of variation of the ratio between measured and standardized pupil size went from 0.03 for our method to 0.09 for Nakayama’s method. We hypothesize that the advantage in our model resulted from the calibration process required to adapt to different subjects and monitors. Furthermore, our exponential model, although more complex and computationally expensive, predicts pupil size as a function of luminosity more accurately. Our model better predicts self-reported arousal starting from pupil size compared to the hyperolic model, going from an average R2 of 0.10 0.087 to an average R2 of 0.436 0.125 (see also Table VIII).
Other linear models, for example, Raiturkar et al. [raiturkar2016decoupling] and Asano et al. [asano2021neural, asano2021pupil] have given even more inferior results with applied to our dataset (see table VIII). Nevertheless, the models of Raiturkar and Asano are particularly interesting because, in addition to developing a calibration procedure, as we did, they have taken into account the temporal response of pupil dilation, something we have not done. Our method relies on average pupil size over relatively long time intervals of at least 5 seconds, making the dynamic aspects that occur over hundreds of milliseconds irrelevant. In fact, we used videos of at least 8 seconds, like most video clips with emotional content. However, it is possible to have real cases in which it is important to measure emotional arousal with stimuli of much shorter duration. Therefore, in future research, we propose to extend our model to account for dynamic changes. We are working to combine our approach with the dynamic approach of Raiturkar et al. [raiturkar2016decoupling] and to further make it automatic rather than requiring the manual adjustment of the time delay for each participant and each video, which introduces potential bias and lacks standardized methodology. An intriguing approach to developing a dynamic model involves employing a neural network incorporating temporal parameters, as demonstrated in [asano2021neural, asano2021pupil].
@@ IV Discussion
In this study, we applied existing methodologies from the literature wherever possible to maximize the value of our dataset, which comprises a relatively large sample of over 40 participants, a collection of over 30 emotional video clips, and self-reported arousal ratings for each participant and each video. To our knowledge, no previous research on pupil size measurement has incorporated self-reported arousal ratings with this level of detail. Some authors have hypothesized and described in their articles the emotional content of the videos shown to the participants without, however, having feedback from the participants themselves([tarnowski2020eye], [nakayama2021controlling], [asano2021neural, asano2021pupil]). Raiturkar and colleagues [raiturkar2016decoupling] used a less subjective approach based on the evaluations of three independent judges and used these as ground truths, generating “excitement scores” for the emotionally involved scenes. However, we contend that this method may not fully capture individuals’ emotional perception variability. We hypothesize that not just physiological responses but also conscious emotional perception and the link between the two vary significantly from individual to individual and are influenced by subjective experiences. To demonstrate this hypothesis, we deliberately disregarded the emotions reported by our 47 participants in the study and instead used the emotions reported by 10 independent judges as ground truth. As reported in paragraph III-C, we trained our model on the pupil size corrected for luminosity in the 47 participants, using the average arousal reported by the 10 independent judges as ground truth. The predictive capacity of our model worsened considerably because the pupil size recorded from the 47 participants could not be significantly linked to the excitement reported by other people (the 10 independent judges). In our case, this is evident because reported arousal levels for a specific video clip varied considerably across participants. Therefore, we believe that it is essential to consider individual differences and cognitive biases in studies of emotional perception. In a future study, we aim to show that the pupils’ responses to emotional stimuli can also vary according to health conditions [grujic2024neurobehavioral], personality traits [bradley2008pupil], and emotional sensitivity [laurenzo2016pupillary].
Our model effectively predicts emotional arousal with high accuracy, surpassing the results of uncorrected pupil size metrics. Its ability to adjust to different luminosity conditions and accommodate individual variations represents a notable enhancement compared to current models. Our novel approach incorporates several innovative features. First, by recognizing that pupillary responses vary across individuals [pamplona2009photorealistic], we ensured that our model accounts for these individual differences, enhancing its generalizability. Second, we addressed the fluctuating nature of ambient luminosity in real-world scenarios, a critical factor often overlooked in controlled laboratory studies. Third, our methodology considers the varying luminosity levels of different visual stimuli, including the inherent variability in monitor brightness and contrast settings. Taken together, we have addressed previous methodological obstacles through a calibration video technique that normalizes these multiple sources of luminosity variation, significantly improving the accuracy of emotion detection.
Our two-stage model accounts for pupil changes due to luminosity and isolates the emotional component from the pupil size measure, offering greater precision and broader application than existing methods. Based on the groundwork laid by previous research, our approach provides a more comprehensive solution for the real-world complexities of experiencing emotions. In the present study, we aimed to demonstrate that the pupil size corrected for luminosity generated by our Python software correlates well with self-reported arousal without needing further processing. This makes our software easy to use, even without advanced knowledge of mathematics, machine learning, or neurobiology. Therefore, we deliberately confined the greater complexity to the first stage of our model, the one responsible for calculating the portion of pupil size due to luminosity. The second stage is very simple and consists of a linear regression that relates arousal to pupil size. Still better results may be obtained by hypothesizing a more complex relationship than a linear one and by applying more complex machine learning and deep learning techniques to the luminosity-corrected pupil size time series. In future studies, we will make our model dynamic, extend it to extreme lighting conditions, such as when wearing eye-tracker glasses outside a laboratory on a sunny day, and adapt it to experiments without eye-trackers and with regular webcams in online experiments.
@@
Due to its ease of implementation, our model has practical applications, particularly in mental health monitoring and consumer research. It requires only existing human insights software, such as iMotions, and basic eye-tracking hardware [imotionsPupillometry101]. No advanced skills are required to detect emotional arousal from pupil size, even without using more complex biosignals to analyze, such as GSR, EEG, and ECG. In the mental health field, where a shortage of trained counselors persists [jonnalagadda2023ensemble], our model offers an accessible AI-driven solution for emotion detection, benefitting under-served populations. In consumer research, pupillometry has been used to assess emotional responses to advertisements but has not seen much popularity among marketers because of its low reliability [Wang2008Validity]. By eliminating luminosity effects, our model enables marketers to gauge the emotional arousal elicited by their advertisements much more accurately. Extensive trials in real-world settings will be essential to improve generalizability, and models should be tested across diverse populations and stimuli.
V Conclusion
In conclusion, our research represents a significant step in emotion arousal detection using physiological indicators. By developing a model that effectively isolates emotional responses from pupillary reactions to luminosity, we have laid the groundwork for more accurate and versatile emotion detection systems. As we continue to refine and expand this technology, the potential for enhancing our understanding of human emotions and their applications in various domains remains exciting and profound.
Acknowledgment
@@
We would like to thank Prof. Luca Ricolfi for inspiring the psychometric methods used, Miss Akashdeep Nijjar, Ms Sophia Carbonero and Dr. Katerina Bourazeri for being an essential member of the research group, Dr. Arno Onken and Miss Cristina Del Prete for the precious advice, Miss Kaylee Shurety, Ms Flavia Popescu-Richardson and Mr. Ville Karhusaari for the continuous, kind and competent assistance, and Miss Domiziana Falaschi for proofreading the manuscript.

### v2

@@ 3.3 Testing our model with ground truth obtained from independent judges rather than being self-reported
3.4 Results of the Gradient Boost Regression-based Self-reported Arousal Prediction Model (GBR-SAPM)
This section presents the results obtained using the GBR-SAPM model. Again, taking into account the effect of luminosity on pupil size had a significant effect. In fact, correcting pupil size for luminosity significantly improved the model’s performance, as shown in Table 9. The model trained on corrected features achieved an of 0.556 0.085, an of 0.229 0.022, and a statistically significant Pearson correlation of 0.765 0.047 (mean p , max p ), across all participants. These results suggest that the corrected pupil size features provided a reliable basis for predicting arousal. In contrast, the model trained on non-corrected features showed much weaker performance across all participants: an of 0.235 0.097, an of 0.301 0.020, and a non-significant Pearson correlation of 0.521 0.116 (mean p = 0.345, max p = 0.863), emphasizing the degradation in model accuracy when luminosity is not accounted for.
| Correction | Target | Score | Pearson (p ) | |
| Corrected | Arousal | 0.556 0.085 | 0.229 0.022 | 0.765 0.047 (mean p ,max p ) |
| Valence | 0.259 0.251 | 0.295 0.041 | 0.595 0.082 (mean p = 0.0012, max p = 0.075) | |
| Non-Corrected | Arousal | 0.235 0.097 | 0.301 0.020 | 0.521 0.116 (mean p = 0.345, max p = 0.863) |
| Valence | 0.205 0.242 | 0.306 0.039 | 0.566 0.065 (mean p = 0.0020, max p = 0.013) |
For valence, as expected, the results were worse and the benefit of the luminosity correction was less pronounced. The corrected model achieved an of 0.259 0.251 and a Pearson correlation of 0.595 0.082 (mean p = 0.0012, max p = 0.075) compared to 0.205 0.242 and 0.566 0.065 (mean p = 0.0020, max p = 0.013) for the non-corrected model, across all participants. The also slightly improved from 0.306 0.039 to 0.295 0.041, as presented in Table 9.
We can conclude that performance improves when moving from a linear regression model (ADM) to a more complex one such as XGBoost (GBR-SAPM). In fact, when the brightness effect is corrected, goes from 0.43 0.12 to 0.556 0.085, NMRSE goes from 0.27 0.036 to 0.229 0.022 and Pearson’s correlation goes from 0.65 0.12 (mean p = 0.0025) to 0.765 0.047 (mean p ).
In Figure 13, we have reported the predicted arousal values compared to the self-reported ones, for both the ADM and GBR-SAPM models and for all participants. The GBR-SAPM model has greater predictive power than the ADM model, and the slope of the regression line of predicted arousal points vs. self-reported arousal goes from 0.22 for the ADM model to 0.89 for the GBR-SAPM model. One reason why the GBR-SAPM model outperforms the ADM model may be that it does not assume a linear relationship between total pupil size, the part driven by emotional arousal, and the part driven by luminosity, which is not true, as recently discoveredPan et al. (2022).
4 Discussion
The aim of this study was to develop and validate a scalable, non-linear computational method for separating the Pupillary Light Response (PLR) from emotional arousal during the viewing of naturalistic videos. Our results demonstrate that the isolated arousal-sensitive residual provides a significantly more accurate representation of subjective emotional states compared to raw pupil diameter. A secondary aim was to create a realistic tool for researchers to be used in typical lab conditions, without the need for specialized equipment.
4.1 Physiological Validity and the Psycho-sensory Residual
Our results provide strong empirical support for the “final common pathway” theory Mathôt (2018); Strauch et al. (2022), which proposes that the pupil acts as an integrated readout where independent neural inputs from both the light reflex and psychological responses are combined Mathôt (2018); Strauch et al. (2022). By effectively suppressing PLR through our calibrated exponential model, we have isolated the underlying autonomic activation, referred to here as the PPR component. The superior performance of both the simple linear regression-based model (ADM) and the more complex model (GBR-SAPM) when using the corrected residual compared to the raw signal suggests that
luminosity-driven fluctuations act as non-linear noise that obscures affective information. Our results do not constitute a direct verification of the “final common pathway” theory, but do demonstrate that, once the variance related to luminosity is eliminated, pupil responses show a stronger and more systematic relationship with subjective self-reported arousal, as would be expected if autonomic activation reflected a convergent pathway. This validation aligns with the decoupling approach proposed by Raiturkar et al. Raiturkar et al. (2016), whilst also providing a more precise physiological fit through non-linear modeling rather than linear approximation. Furthermore, the technical rigor of our exponential fit is supported by advanced optical studies, such as the Spatially Weighted Corneal Flux Density approach by Zhang et al. Zhang et al. (2019), even though our model simplifies these principles for laboratory usability.
4.2 Two-Stage Modeling
A goal of this research was to ensure that the luminosity-corrected signal is immediately usable for affective inference without requiring exhaustive post-processing. To this end, we developed two distinct modeling stages. The first stage, the Arousal Detection Model (ADM), utilizes a simple linear regression to relate the corrected pupil residual to arousal. This was a deliberate design choice to demonstrate that once the complex non-linear effects of luminosity are removed, even a straightforward linear approach yields significant predictive improvements. This makes our methodology highly accessible to researchers in psychology and neurobiology who may not have specialized expertise in advanced machine learning. In the second stage, we implemented the more sophisticated Gradient Boosted Regression-based Self-reported Arousal Prediction Model (GBR-SAPM). We achieved superior performance metrics, illustrating that, although the corrected signal is robust in its raw form, advanced machine learning can further resolve intricate physiological dynamics.
4.3 Multiplicative Light-Arousal Interaction and Limit of the Additive Model
A critical finding in this work is that while linear models often perform well under limited luminosity variations, they fail to capture the inherently non-linear and multiplicative interaction between arousal and light. As demonstrated by Pan et al. Pan et al. (2022), emotional arousal effects are not monolithic and interact differently depending on background luminance, with strongest effects occurring at very low luminances. Our model accounts for this non-linear structure and provides a more precise physiological fit than linear approximations. This distinction is vital for laboratory realism, where screen-based stimuli vary unpredictably. By utilizing an exponential fit (LEPM) and a non-linear regressor (GBR-SAPM), we move beyond the “additive” assumption that characterizes simpler decoupling frameworks.
4.4 Methodological Scalability and Ecological Validity
A significant innovation of this work is its scalability for typical psychological research. Traditional experimental designs require the use of controlled visual stimuli and isoluminant conditions to achieve stable and uniform luminosity across stimuli, an artificial process that is laborious and difficult to adapt to videos. Our method eliminates this requirement with a one-time individual calibration procedure that accounts for inter-subject differences and monitor configurations Pamplona et al. (2009). We have also taken into account the fluctuating nature of ambient light, a critical factor that is often overlooked in controlled laboratory studies. Our work will allow researchers to utilize diverse video content without specialized hardware like photometers, bridging the gap between high-precision engineering models and the practical operational needs of behavioral science.
4.5 Holistic Experience vs. Dual-Task Confound
Our research focused specifically on the holistic emotional response to short video clips. We intentionally used brief stimuli to evoke a single, consistent emotional state, and our ground truth was obtained through a single post-stimulus self-report. Our stimuli were designed to elicit a single, relatively stable emotional state per clip, and because clip durations (at least 8 seconds) exceeded the typical latency of pupillary responses, we considered the average value of pupil size over the entire video rather than transient dynamics. This choice prioritizes robustness and interpretability for holistic emotional states, though we acknowledge that dynamic features may carry additional information in paradigms involving rapid affective fluctuations. Although greater temporal granularity is a common computational goal, requiring participants to provide continuous, second-by-second self-reports can introduce dual-task interference. In such scenarios, the cognitive effort of monitoring and reporting one’s internal state can compromise the authenticity of the emotional experience. However, there are situations, such as when playing a video game, where it makes sense to provide continuous feedback on one’s state of arousal, using special devices, like RankTrace Lopes et al. (2017).
4.6 Validation Strategy: Self-Report vs. Independent Judges
@@ 4.4 Methodological Scalability and Ecological Validity
In this study, we used a relatively large sample, compared to existing research, with over 40 participants, a collection of over 30 emotional video clips, and self-reported arousal ratings for each participant and each video. To our knowledge, no previous research on pupil size measurement has incorporated self-reported arousal ratings with this level of detail. Some authors have hypothesized and described in their articles the emotional content of the videos shown to participants without validation from the participants themselves Tarnowski et al. (2020), Nakayama et al. (2021), Asano et al. (2021a, b). Raiturkar and colleagues Raiturkar et al. (2016) used a less subjective approach based on the assessments of three independent judges and used them as ground truth, generating “arousal scores” for emotionally engaging scenes. A key finding in our study, however, was the need to use participants’ self-assessments as the ground truth for interpreting biosignals. Indeed, when we used the ratings of 10 independent judges as a counter-control, the predictive accuracy of the model deteriorated significantly, with a 188% increase in . This highlights that the pupillary response is an idiosyncratic physiological indicator, influenced by individual subjective experiences and cognitive biases. Although independent judges may be useful for the normative classification of stimuli, self-assessment is scientifically more consistent for training machines to interpret an individual’s specific emotional state. Our study shows that physiological responses, conscious emotional perception and their combination vary significantly between individuals. In our study, this is evident because the arousal and valence reported for a given video clip varied greatly between participants, underscoring the need to account for individual differences in studies of emotional perception.
4.7 Practical Applications and Accessibility
@@
Thanks to its ease of implementation, our method has important practical applications, particularly in mental health monitoring and neuro-marketing. It requires only standard software for analyzing human perceptions, such as iMotions, and basic hardware for eye tracking iMotions , making it a cost-effective solution for myriad research contexts. Our method allows for the detection of emotional arousal without the need to analyze more complex or invasive biosignals, such as GSR, EEG, or ECG. In the field of mental health, where there is a persistent shortage of qualified counsellors, this approach offers an accessible, AI-based solution for emotion detection that could benefit under-served populations Jonnalagadda et al. (2023). In consumer research, pupillometry has been used to assess emotional responses to advertisements but has not seen much popularity among marketers because of its low reliability Wang and Minor (2008). By eliminating luminosity effects, our model enables marketers to gauge the emotional arousal elicited by their advertisements much more accurately. Although extensive testing in real-world contexts is still needed to improve generalizability across different populations and stimuli, our method provides a solid foundation for more reliable affective assessment in these fields.
4.8 Limitations and Future Directions
In future work, we will include localized ocular factors in our model, such as gaze position or spatial light weighting, as discussed in the technical literature. Some of these factors, such as gaze position, are implicitly taken into account but not used as features. A limitation of our work is that our models lack of granularity. In future work, we will make our model dynamic and increase temporal granularity without increasing the burden on participants. For this purpose, we want to use deep learning techniques, such as a neural network that incorporates temporal parameters, as demonstrated in Asano et al. (2021a, b). While we believe the current GBR-SAPM model accounts for the multiplicative light-arousal interaction implicitly through its non-linear architecture, future developments will focus on accounting for this phenomenon explicitly to further refine the separation of signals. Furthermore, while this article has focused exclusively on the pupil to isolate its predictive value, future work will integrate this signal with multimodal data, including heart rate and GSR, which were recorded during this study. In a future study, we will also explore the extent to which pupil responses to emotional stimuli also vary depending on health conditions Grujic et al. (2024), personality traits Bradley et al. (2008), and emotional sensitivity Laurenzo et al. (2016).
5 Conclusion
This research represents a significant advancement in the use of pupillometry for emotional arousal detection in naturalistic laboratory settings. By developing a model that effectively isolates the psychosensory component from the dominant pupillary light reflex, we have provided a solution to a long-standing methodological obstacle in affective computing.
Our findings establish three critical conclusions:
A) Correction is Mandatory: Raw pupil size lacks significant predictive power for measuring arousal during dynamic video viewing; accurate inference is only possible after removing the non-linear luminosity component.
B) Subject-Centric Validation: For the purposes of training machines in which the target variable is the arousal experienced by participants, participants’ self-reports provide a more appropriate reference than external judges’ assessments, which reflect the normative properties of the stimulus rather than subjective experience.
C) Methodological Scalability: The proposed calibrated exponential model offers a “realistic and scalable” alternative to complex engineering approaches, enabling high-fidelity research with standard eye-tracking hardware and diverse video stimuli.
While the current work focuses on holistic emotional responses to short video clips to preserve ecological validity and avoid dual-task reporting confounds, future research will extend this framework into a dynamic model. By integrating this corrected pupillary signal with other physiological measures such as GSR and heart rate, we aim to develop a robust, multi-modal system for real-time emotion detection across various domains, including healthcare and human-computer interaction.
Acknowledgements.
@@
We would like to thank Prof. Luca Ricolfi for inspiring the psychometric methods used, Akashdeep Nijjar, Sophia Carbonero and Dr Katerina Bourazeri for being an essential member of the research group, Dr Arno Onken and Cristina Del Prete for the precious advice, Kaylee Shurety, Flavia Popescu-Richardson and Ville Karhusaari for the continuous, kind and competent assistance, and Domiziana Falaschi for proofreading the manuscript.References
- A neural-network-based eye pupil reaction model for use with television programs. Pupil Reactions in Response to Human Mental Activity, pp. 31–47. Cited by: §1, Table 8, §4.6, §4.8.
- Pupil reaction model using a neural network for brightness change. Pupil Reactions in Response to Human Mental Activity, pp. 15–30. Cited by: §4.6, §4.8.
- Measuring emotions in education using wearable devices: a systematic review. Computers & Education 200, pp. 104797. Cited by: §1.
- Facial emotion recognition through artificial intelligence. Frontiers in Computer Science 6, pp. 1359471. Cited by: §1.
- The pupil as a measure of emotional arousal and autonomic activation. Psychophysiology 45 (4), pp. 602–607. Cited by: §1, 1st item, §4.8.
- The impact of emotion on perception, attention, memory, and decision-making. Swiss medical weekly 143 (1920), pp. w13786–w13786. Cited by: §1.
- Handbook of psychophysiology. Cambridge university press. Cited by: 2nd item.
- Open-dpsm: an open-source toolkit for modeling pupil size changes to dynamic visual inputs. Behavior Research Methods 56 (6), pp. 5605–5621. Cited by: §1.
- Background luminance effects on pupil size associated with emotion and saccade preparation. Scientific reports 10 (1), pp. 15718. Cited by: §1, §1.
- Facial emotion recognition in children with externalising behaviours: a systematic review. Clinical child psychology and psychiatry 25 (4), pp. 1068–1085. Cited by: §1.
- Alexithymia and emotional processing: a longitudinal mixed methods research. Research in Psychotherapy: Psychopathology, Process, and Outcome 21 (1), pp. 292. Cited by: §2.2.1.
- Emotion recognition for healthcare surveillance systems using neural networks: a survey. In 2021 International Wireless Communications and Mobile Computing (IWCMC), pp. 681–687. Cited by: §1.
- Attenuation of the dynamic pupil light response during screen viewing for arousal assessment. Frontiers in virtual reality 3, pp. 971613. Cited by: §1, §2.1.3.
- From pre-processing to advanced dynamic modeling of pupil data. Behavior Research Methods 56, pp. 2234–2257. External Links: Document Cited by: §1.
- Monitoring and processing of the pupil diameter signal for affective assessment of a computer user. In International Conference on Human-Computer Interaction, pp. 49–58. Cited by: §1.
- Review on psychological stress detection using biosignals. IEEE transactions on affective computing 13 (1), pp. 440–460. Cited by: §1.
- Emotion processing in depression and anxiety disorders in older adults: systematic review. BJPsych Open 7 (1), pp. e7. Cited by: §2.2.1.
- Emotion regulation: current status and future prospects. Psychological inquiry 26 (1), pp. 1–26. Cited by: §1.
- Neurobehavioral meaning of pupil size. Neuron 112 (20), pp. 3381–3395. Cited by: §4.8.
- Autonomic nervous system disorders. In Hankey’s Clinical Neurology, pp. 749–756. Cited by: §1.
- Pupil size estimation model based on luminance and chromaticity spatial distribution. Journal of the Society for Information Display 31 (4), pp. 184–194. Cited by: §1.
- [22] Pupillometry 101: how pupil responses reveal brain activity. Note: Accessed: February 12, 2025 External Links: Link Cited by: §2.2.1, §4.7.
- Emotion theory and research: highlights, unanswered questions, and emerging issues. Annual review of psychology 60 (1), pp. 1–25. Cited by: §1.
- An ensemble-based machine learning model for emotion and mental health detection. Journal of Information & Knowledge Management 22 (02), pp. 2250075. Cited by: §4.7.
- The role of emotional intelligence in leadership effectiveness. Scholar’s Digest: Journal of Commerce & Management 1 (1), pp. 84–93. Cited by: §1.
- The effects of measuring emotion: physiological reactions to emotional situations depend on whether someone is asking. PloS one 8 (6), pp. e64959. Cited by: §1.
- Identifying core affect in individuals from fmri responses to dynamic naturalistic audiovisual stimuli. PloS one 11 (9), pp. e0161589. Cited by: §2.2.1, §2.2.1, §2.2.2.
- A solid frame for the window on cognition: modeling event-related pupil responses. Journal of vision 16 (3), pp. 28–28. Cited by: §1.
- Emotion processing deficits: a liability spectrum providing insight into comorbidity of mental disorders. Neuroscience & Biobehavioral Reviews 52, pp. 153–171. Cited by: §2.2.1.
- Preprocessing pupil size data: guidelines and code. Behavior Research Methods 51 (3), pp. 1336–1342. External Links: Document Cited by: §1.
- Pupillary response abnormalities in depressive disorders. Psychiatry research 246, pp. 492–499. Cited by: §4.8.
- Empirical investigation of multi-source cross-validation in clinical machine learning. arXiv preprint arXiv:2403.15012. Cited by: item 1.
- Emotion and decision making. Annual review of psychology 66 (1), pp. 799–823. Cited by: §1.
- Ranktrace: relative and unbounded affect annotation. In 2017 Seventh International Conference on Affective Computing and Intelligent Interaction (ACII), pp. 158–163. Cited by: §4.5.
- Role of sympathetic and parasympathetic systems in reflex dilatation of the pupil: pupillographic studies. Archives of Neurology & Psychiatry 64 (3), pp. 313–340. Cited by: §1.
- Methods in cognitive pupillometry: design, preprocessing, and statistical analysis. Behavior Research Methods 54, pp. 3055–3077. External Links: Document Cited by: §1.
- Pupillometry: psychology, physiology, and function. Journal of Cognition 1 (1). Cited by: §1, §4.1.
- Influences of luminance and accommodation stimuli on pupil size and pupil center location. Investigative ophthalmology & visual science 55 (4), pp. 2166–2172. Cited by: §1.
- Temporal alignment of pupillary response with stimulus events via deconvolution. The Journal of the Acoustical Society of America 139 (3), pp. EL57–EL62. Cited by: §1.
- Controlling the effects of brightness on the measurement of pupil size as a means of evaluating mental activity. Pupil Reactions in Response to Human Mental Activity, pp. 1–14. Cited by: §1, Table 8, §4.6.
- Modeling human pupil dilation to decouple the pupillary light reflex. Open Physics 17 (1), pp. 458–467. Cited by: §1.
- Photorealistic models for pupil light reflex and iridal pattern deformation. ACM Transactions on Graphics (TOG) 28 (4), pp. 1–12. Cited by: §1, §2.1.3, §4.4.
- Arousal-based pupil modulation is dictated by luminance. Scientific reports 12 (1), pp. 1390. Cited by: §1, §2.3, §3.4, §4.3.
- The effects of emotional arousal on pupil size depend on luminance. Scientific Reports 14 (1), pp. 21895. Cited by: §1, §1, §1.
- Towards an accurate measure of emotional pupil dilation responses: a model for removing the effect of luminosity. In 2024 IEEE International Conference on Metrology for eXtended Reality, Artificial Intelligence and Neural Engineering (MetroXRAINE), pp. 1–6. Cited by: item 1, §1, §2.1.1, §2.1.1, §2.1.2, §2.1.3, §2.1.3, §2.1.3, §2.1.4.
- Pupil size variation as an indication of affective processing. International journal of human-computer studies 59 (1-2), pp. 185–198. Cited by: §1.
- Emotion in human-computer interaction. Expanding the frontiers of visual analytics and visualization, pp. 239–262. Cited by: §1.
- Decoupling light reflex from pupillary dilation to measure emotional arousal in videos. In Proceedings of the ACM Symposium on Applied Perception, pp. 89–96. Cited by: §1, §1, §3.3, Table 8, Table 8, §4.1, §4.6.
- Multimodal emotion recognition: a comprehensive review, trends, and challenges. Wiley Interdisciplinary Reviews: Data Mining and Knowledge Discovery 14 (6), pp. e1563. Cited by: §2.2.1.
- A circumplex model of affect.. Journal of personality and social psychology 39 (6), pp. 1161. Cited by: §2.2.1, §2.2.1.
- A review of emotion recognition using physiological signals. Sensors 18 (7), pp. 2074. Cited by: §1.
- Pupillometry. Wiley Interdisciplinary Reviews: Cognitive Science 5 (6), pp. 679–692. Cited by: §1.
- Multimodal emotion recognition in response to videos. IEEE Transactions on Affective Computing 3 (2), pp. 211–223. External Links: Document Cited by: §1.
- The pupils. Clinical Methods: The History, Physical, and Laboratory Examinations. 3rd edition. Cited by: §1.
- Pupillometry as an integrated readout of distinct attentional networks. Trends in Neurosciences 45 (8), pp. 635–647. Cited by: §1, §4.1.
- Emotion recognition of the singing voice: toward a real-time analysis tool for singers. arXiv preprint arXiv:2105.00173. Cited by: §1.
- Eye-tracking analysis for emotion recognition. Computational intelligence and neuroscience 2020. Cited by: §1, §2.3, §4.6.
- The influences of emotion on learning and memory. Frontiers in psychology 8, pp. 235933. Cited by: §1.
- How emotions regulate social life: the emotions as social information (easi) model. Current directions in psychological science 18 (3), pp. 184–188. Cited by: §1.
- Integrating artificial intelligence to assess emotions in learning environments: a systematic literature review. Frontiers in Psychology 15, pp. 1387089. Cited by: §1.
- Validity, reliability, and applicability of psychophysiological techniques in marketing research. Psychology & Marketing 25 (2), pp. 197–232. External Links: Document, Link Cited by: §4.7.
- Emotion recognition based on multimodal physiological electrical signals. Frontiers in Neuroscience 19, pp. 1512799. Cited by: §1.
- A unified formula for light-adapted pupil size. Journal of vision 12 (10), pp. 12–12. Cited by: §1, §2.1.3.
- Pupil dilation deconvolution reveals the dynamics of attention at high temporal resolution. Proceedings of the National Academy of Sciences 109 (22), pp. 8456–8460. Cited by: §1.
- Ocular autonomic nervous system: an update from anatomy to physiological functions. Vision 6 (1), pp. 6. Cited by: §1.
- Development of emotional intelligence: towards a multi-level investment model. Human development 46 (2-3), pp. 69–96. Cited by: §1.
- Emotion recognition using multi-modal data and machine learning techniques: a tutorial and review. Information Fusion 59, pp. 103–126. Cited by: §2.2.1.
- Pupil size estimation based on spatially weighted corneal flux density. IEEE Photonics Journal 11 (6), pp. 1–9. Cited by: §1, §2.1.3, §2.1.3, §4.1.